OpenStudy (anonymous):

I. DONT. UNDERSTAND. THIS!! ='(Solve: x^2 - 20x + 36 = 0 {6, -6} {2, 18} {-4, -9} no solution

OpenStudy (anonymous):

Hello. You can factor the equation as \((x - 2)(x - 18) = 0\) or you can use quadratic formula: \[ x_{1,2} = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}. \]